OpenStudy (anonymous):

Which ordered pair is a solution to the equation 4x + 7y = –5? A. (–4, –3) B. (–4, 3) C. (4, –3) D. (4, 3)

OpenStudy (tfht):

get it so that y=mx+b first

OpenStudy (tfht):

so you would get 7y=4x-5 then divide both sides by 7 to get y=(4/7)x-(5/7)

OpenStudy (anonymous):

so the answers b?

OpenStudy (mathstudent55):

There's no need to do anything to the equation. Just plug in each pair into the equation and see which ordered pair makes the equation true.

OpenStudy (anonymous):

i'm having a hard time choosing between b and c

OpenStudy (mathstudent55):

4x + 7y = -5 Try A. (-4, -3) Is 4(-4) + 7(-3) equal to -5? -16 - 21 = -37. -37 is not equal to -5, so choice A is not it. Now do the same with the other choices until one of them works.

OpenStudy (anonymous):

alright

OpenStudy (tfht):

sorry, i had to leave for a short time.... but if you did get in y=mx+b form, you could just start at (-5/7) then go up 4 and to the right 7

OpenStudy (tfht):

but yeah the other way is easier...

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Joey fixes bicycles and charges a shop fee of $15 and $10 per hour for time worked. If t is the number of hours worked and C is the total charge for completed work, then the variables are related by the equation C = 10t + 15. Which variable is the independent variable? A. There is no relationship between the variables. B. C C. t

OpenStudy (tfht):

c

OpenStudy (tfht):

an independent variable is a variable which can be assigned any permissible value without any restriction imposed by any other variable.

OpenStudy (anonymous):

thank you very much :)

OpenStudy (tfht):

or you could just say that it is the variable in the equation. and no problem!

OpenStudy (anonymous):

you're awesome :D

OpenStudy (tfht):

lol thanks

OpenStudy (anonymous):

:)